OpenStudy (anonymous):

Solve the following systems of equations using the graphing method. What type of system is it? Name the solution if there is one. . 3x-y=-2 3x-y=4 show graph

TheSmartOne (thesmartone):

You can use an online graphing calculator such as Desmos Graphing Calculator

OpenStudy (anonymous):

? explain please

OpenStudy (anonymous):

Slope-intercept form, then graph

OpenStudy (anonymous):

3x-y=-2 /3 /3 x-y=-2/3 +y +y x=-2/3y 3x-y=4 -3x -3x -y=4-3x /- /- y=-1x

OpenStudy (anonymous):

@megan_tanuis

OpenStudy (anonymous):

@megan_tanuis

OpenStudy (anonymous):

You need to subtract 3x from both sides

OpenStudy (anonymous):

@megan_tanus where do i subtract 3x from? I'm confused

OpenStudy (anonymous):

@megan_tanuis

TheSmartOne (thesmartone):